### v2

# Dark matter in QCD-like theories with a theta vacuum: cosmological and astrophysical implications

Abstract
QCD-like theories in which the dark matter (DM) of the Universe is hypothesized to be a thermal relic in the form of a dark pion has been extensively investigated, with most studies neglecting the CP-violating $\theta$-angle associated with the topological vacuum. We point out that a non-vanishing $\theta$ could potentially trigger resonant number-changing processes giving rise to the observed relic density in agreement with perturbative unitarity as well as observations of clusters of galaxies. This constitutes a novel production mechanism of MeV DM and an alternative to those relying on the Wess-Zumino-Witten term. Moreover, for specific meson mass spectra, similar resonant scatterings serve as a realization of velocity-dependent self-interacting DM without a light mediator. Explicit benchmark models are presented together with a discussion of possible signals, including gravitational waves from the chiral phase transition associated with the dark pions.
Early studies of the strong-CP problem of the Standard Model (SM) Shifman:1979if ; Crewther:1979pi found that the topological $\theta$-vacuum in QCD induces cubic interactions among the pseudo-Goldstone bosons of the meson octet, which are absent for $\theta=0$. In particular, the SM $\eta$ meson may decay into two pions with a rate proportional to $\theta^{2}$. Similar interactions for pion dark matter (DM) in QCD-like theories naturally induce number-changing processes in the Early Universe (see Fig. 1), which alter the DM amount that goes out of equilibrium from the primordial plasma. We will show that this constitutes a novel production mechanism of DM in QCD-like theories. This provides an alternative to the popular SIMP model Hochberg:2014kqa of DM at the MeV scale based on $3\to 2$ annihilations Carlson:1992fn ; Hochberg:2014dra induced by a topological 5-point interaction, the so-called Wess-Zumino-Witten (WZW) term Wess:1971yu ; Witten:1983tw . We refer the reader to Hochberg:2015vrg ; Kuflik:2015isi ; Bernal:2015bla ; Bernal:2015xba ; Bernal:2015ova ; Kuflik:2015isi ; Choi:2015bya ; Choi:2016hid ; Soni:2016gzf ; Kamada:2016ois ; Bernal:2017mqb ; Cline:2017tka ; Choi:2017mkk ; Kuflik:2017iqs ; Heikinheimo:2018esa ; Choi:2018iit ; Hochberg:2018rjs ; Bernal:2019uqr ; Choi:2019zeb ; Katz:2020ywn ; Smirnov:2020zwf ; Xing:2021pkb ; Braat:2023fhn ; Bernreuther:2023kcg ; Garani:2021zrr ; Dey:2016qgf for related work on MeV DM, and to  Redi:2016kip ; Draper:2018tmh ; Abe:2024mwa for studies of the effect of the $\theta$ angle in QCD-like DM models in other contexts.
Likewise, in the presence of a non-vanishing $\theta$, resonances similar to the $\eta$ meson may mediate velocity-dependent scatterings among DM pions, see Fig. 1. Such resonances have astrophysical implications Chu:2018fzy , as velocity-dependent scatterings can diminish the central density of DM halos hosting galaxies Spergel:1999mh ; Tulin:2017ara , while circumventing the stringent constraints on DM self-interactions derived from observations of galaxy clusters Harvey:2015hha ; Bondarenko:2017rfu ; Harvey:2018uwf ; Sagunski:2020spe ; DES:2023bzs .
We will illustrate this in two specific benchmark models, one of them resembling SM QCD. We structure the discussion as follows: the first section highlights the often-overlooked interactions induced by the $\theta$ angle in QCD-like models of DM, whereas the following two sections present ensuing cosmological and astrophysical consequences. Finally, we will present a summary along with a potential outlook. In the Appendix we provide the details of our calculations.
Pion DM in a $\theta$-vacuum. Although our conclusions apply to other gauge groups, we focus on a $SU(N_{c})$ QCD-like theory of DM, with $N_{c}\geq 3$, and $N_{f}$ flavors of light Dirac quarks in the fundamental representation. As usual, the Lagrangian reads
[…]
Figure 1: The $\theta$-vacuum could induce DM number changing processes in the early Universe and velocity-dependent self-scattering in DM halos today.
We assume $\theta\ll 1$ throughout for simplicity, and take $M=\text{diag}(m_{1},\cdots,m_{N_{f}})$, with $m_{q}\neq 0$ for all flavors. Due to the chiral anomaly, under an arbitrary transformation $q_{L,R}\to e^{\mp i\theta Q/2}q_{L,R}$, the angle in Eq. (1) shifts as $\theta\to\theta(1-{\rm Tr}\,Q)$ while $M\to M_{\theta}=e^{i\theta Q/2}Me^{i\theta Q/2}$. Taking a transformation with ${\rm Tr}\,Q=1$, we move the $\theta$ parameter from the $F\widetilde{F}$ term to the quark mass matrix .
In analogy with ordinary QCD, we expect that strong interactions confine at some energy scale $\Lambda$, giving rise to a fermion condensate, $\langle\bar{q}q\rangle\sim\Lambda^{3}$. This induces a chiral phase transition (PT) for $m_{q}\ll\Lambda$, which spontaneously breaks the chiral flavor symmetry group of Eq. (1): $SU(N_{f})_{L}\times SU(N_{f})_{R}\to SU(N_{f})_{V}$. This leads to $N_{f}^{2}-1$ pseudo-Goldstone bosons, $\pi^{a}$, described by chiral perturbation theory (see e.g. Refs. Pich:1991fq ; Scherer:2002tk ), according to
$\mathcal{L}_{\rm eff}=\frac{f_{\pi}^{2}}{4}{\rm Tr}[\partial_{\mu}U^{\dagger}\partial^{\mu}U]+\frac{f_{\pi}^{2}}{2}B_{0}{\rm Tr}[M_{\theta}U+U^{\dagger}M_{\theta}^{\dagger}],$ (2)
where $f_{\pi}$ is the dark meson constant, such that $\Lambda\sim 4\pi f_{\pi}/\sqrt{N_{c}}$, and $U=\exp({i\pi^{a}\lambda^{a}/f_{\pi}})$ with $\lambda^{a}$ being the generators of $SU(N_{f})$. Here the condensate is parametrized as $\langle\bar{q}q\rangle\equiv-B_{0}f_{\pi}^{2}$.
The choice $Q=M^{-1}/{\rm Tr}M^{-1}$ yields ${\rm Tr}Q=1$ and no linear terms in $\pi^{a}$ in the effective Lagrangian at leading order in $\theta$ Crewther:1979pi ; DiLuzio:2020wdo
$\!\!\!\mathcal{L}_{\rm eff}\!\supset\frac{f_{\pi}^{2}B_{0}}{2}\left({\rm Tr}[M(U\!+\!U^{\dagger})]\!+\!\frac{i\theta}{{\rm Tr}M^{-1}}{\rm Tr}[U\!-\!U^{\dagger}]\right)\,.$ (3)
The first term includes ordinary interactions involving an even number of mesons such as $\mathcal{L}_{\rm mass}=-B_{0}\pi^{a}\pi^{b}{\rm Tr}[M\{\lambda^{a},\lambda^{b}\}]/4$, to be corrected by ${\cal O}(\theta^{2})$ contributions, while the second term –often neglected– gives rise to interactions with an odd number of mesons [^1] Throughout ${\rm Tr}[\lambda_{a}\lambda_{b}]=2\delta_{ab}$, while the symmetric tensors are $d_{abc}=\frac{1}{4}{\rm Tr}(\{\lambda_{a},\lambda_{b}\}\lambda_{c})$ , and $c_{abcde}=(\delta_{ab}d_{cde}+\delta_{cd}d_{abe})/N_{f}+\frac{1}{2}d_{abm}d_{cdn}d_{mne}$.
$\!\!\!\mathcal{L}_{\theta}=\frac{B_{0}\theta}{3f_{\pi}{\rm Tr}M^{-1}}\bigg{(}d_{abc}\pi_{a}\pi_{b}\pi_{c}-\frac{c_{abcde}}{10f_{\pi}^{2}}\pi_{a}\pi_{b}\pi_{c}\pi_{d}\pi_{e}\bigg{)}.$ (4)
Non-vanishing for $N_{f}\geq 3$, this allows processes absent when $\theta=0$, particularly $s$-wave 3-to-2 annihilations, as pointed out in Ref. Kamada:2017tsq for a degenerate meson spectrum. This may be compared to the topological WZW term, $\mathcal{S}_{\mathrm{WZW}}=-iN_{c}\int\operatorname{Tr}\left(U^{\dagger}\mathrm{d}U\right)^{5}/240\pi^{2}$, which induces a 5-point interaction Wess:1971yu ; Witten:1983tw leading to velocity suppressed 3-to-2 annihilations Hochberg:2014kqa ; Kamada:2022zwb .
For illustration, we consider the benchmark scenarios in table 1. BM1 resembles ordinary QCD, with eight dark mesons referred to as in the SM. After accounting for the mixing angle, $\theta_{\eta\pi}$, the heaviest and lightest are respectively $\eta$ and $\pi^{0}$, with $m_{\pi^{0}}\approx m_{\pi^{\pm}}<m_{K}<m_{\eta}$. On the other hand, BM2 has a remnant $SU(n)$ symmetry under which the $N_{f}^{2}-1$ mesons organize as $n^{2}-1$ scalars in the adjoint, $2n$ scalars in the (anti)fundamental representations, $K/\bar{K}$, and one singlet, the $\eta$ resonance. Neglecting $\mathcal{O}(\theta^{2})$ corrections, the masses squared are respectively $2B_{0}m,\,B_{0}(m+\mu)$ and $2B_{0}\left(m+n\mu\right)/(n+1)$, with $\mu>m$.
Benchmark BM1 BM2
$N_{f}$ 3 $n+1$
$M$ ${\rm diag}(m_{u},m_{d},m_{s})$ ${\rm diag}(m,..m,\mu)$
Spectrum $\pi^{0},\pi^{\pm},K^{0},\bar{K}^{0},K^{\pm},\eta$ $\pi,K,\bar{K},\eta$
DM particle $\pi^{0}\sim$ singlet $\pi\sim$ adjoint of $SU(n)$
$\xi$ in Eq. (5) $\cos^{2}3\theta_{\eta\pi}$ $6(n-1)/n$
$\langle\sigma_{\eta\pi}v\rangle$ in Eq. (7) $\frac{445\sqrt{5}m_{\rm DM}^{2}\delta}{5184\pi f_{\pi}^{4}}$ $\frac{\sqrt{5}m_{\rm DM}^{2}(n^{2}-4)}{192\pi f_{\pi}^{4}n^{2}(n+1)}$
$\sigma_{0}$ in Eq. (8) $\frac{m_{\rm DM}^{2}}{128\pi f_{\pi}^{4}}$ $\frac{m_{\rm DM}^{2}}{64\pi f_{\pi}^{4}}\frac{3n^{4}-2n^{2}+6}{n^{2}(n^{2}-1)}$
Table 1: Benchmark models considered in this work.
Scaling as $m_{\pi}\sim\sqrt{m_{q}\Lambda}$, meson masses are naturally close to each other making plausible a spectrum with resonances. Dark baryons and dark glue-balls have typical masses of order $\Lambda$. This justifies the assumptions that the DM is made of the lightest dark meson. All mesons are stable except $\eta$, as the cubic interaction in Eq. (4) predicts
$\!\!\!\!\!\!\Gamma\left(\eta\rightarrow\!{\rm DM}\,{\rm DM}\right)=\frac{\theta^{2}B_{0}^{2}\xi}{24\pi f_{\pi}^{2}m_{\eta}({\rm Tr}M^{-1})^{2}}\sqrt{1-\frac{4m_{\rm DM}^{2}}{m_{\eta}^{2}}}\,,$ (5)
in agreement with Refs. Shifman:1979if ; Crewther:1979pi ; Pich:1991fq . We introduce
$v_{R}=2\sqrt{\frac{m_{\eta}-2m_{\rm DM}}{m_{\rm DM}}}\,,$ (6)
to study the non-relativistic resonant effects of Fig. 1, which appear for $v_{R}\lesssim 1$. For BM1, up to $\mathcal{O}(v_{R}^{2})$ and $\mathcal{O}(\theta^{2})$ corrections, this fixes $\delta\equiv m_{\pi^{\pm}}/m_{\pi^{0}}-1$ in between 0 and 0.075, depending on $r_{ud}\equiv m_{u}/m_{d}$. Further details on the mass spectrum are in the Appendix.
Refer to caption Refer to caption
Figure 2: Left: the cosmological DM relic abundance, $\Omega_{\rm DM}h^{2}=0.12$ Planck:2018vyg , is reproduced along the plane for different values of $r_{ud}$ in BM1. In the upper gray region DM is overabundant for any physical value of $r_{ud}$, while the blue region is excluded by observations of galaxy clusters, $\sigma_{0}/m_{\rm DM}>0.5\text{ cm}^{2}/\text{g}$. The star corresponds to the benchmark points plotted on the right panel. In the dark-gray region chiral perturbation theory breaks down, $m_{\rm DM}/f_{\pi}>4\pi/\sqrt{N_{c}}$ Kamada:2022zwb . Right: Velocity dependence of the DM self-scattering cross section per unit of mass, for representative values of $v_{R}$ in Eq. (6).
Cosmological implications. The $\theta$ angle impacts DM production due to the number-changing processes resulting from Eq. (4). These include 3-to-2 annihilations of DM and inverse decays ${\rm DM}\,{\rm DM}\to\eta$. As discussed below, in the absence of the resonance, the former process in our benchmarks only leads to the observed relic density in parameter space already excluded observationally. For the following discussion, we assume that the dark sector remains in kinetic equilibrium with the SM, thereby sharing the same temperature $T$. The model-building requirements for this are similar to those of ordinary SIMPs Hochberg:2014kqa .
As in usual scenarios of co-annihilating DM, all the stable mesons eventually convert to DM with a contribution to the DM relic density that depends on their mass. Accordingly, DM and its coannihilating partners are referred to as $\pi_{\rm DM}$. Kaons in both benchmarks are sufficiently heavy so that their contribution is exponentially suppressed and thus negligible. Consequently, $n_{\pi_{\rm DM}}\simeq n_{\pi}$ in BM2. In contrast, due to $\delta\lesssim 0.075$, $\pi^{\pm}$ in BM1 give a non-negligible contribution to the relic density, which can be calculated by noting that fast reactions $\pi^{+}\pi^{-}\leftrightarrow\pi^{0}\pi^{0}$ establish chemical equilibrium, $n_{\pi^{0}}/n_{\pi^{\pm}}=(n_{\pi^{0}}/n_{\pi^{\pm}})_{\rm eq}$, allowing to write a single Boltzmann equation for the combination $n_{\pi_{\rm DM}}=n_{\pi^{0}}+2n_{\pi^{\pm}}$. See Ref. Katz:2020ywn for an extended discussion.
On the other hand, $\eta$ does not behave as a co-annihilating partner but as a catalyzer Chu:2024rrv , since it induces number-changing processes such as $\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$, $\eta\to\pi_{\rm DM}\pi_{\rm DM}$ and their inverse. In light of this, only Boltzmann equations for $\eta$ and $\pi_{\rm DM}$ are necessary.
They take a simple form in the case of most interest for this work, namely when $v_{R}\lesssim 0.1$ and $\theta$ is sufficiently large so that decays and inverse decays, $\eta\leftrightarrow\pi_{\rm DM}\pi_{\rm DM}$, are both active even after all other DM number changing interactions have frozen-out. Then, the process $\pi_{\rm DM}\pi_{\rm DM}\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$ is dominated by the exchange of an on-shell $\eta$ resonance (see Fig. 1). In practice, this means that the 3-to-2 annihilation is a two-step process: $\pi_{\rm DM}\pi_{\rm DM}\to\eta$ followed by $\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$, which justifies including only these in the Boltzmann equations and neglecting the small non-resonant piece Kolb:1979qa .
Noting that (inverse) decays do not affect $n=n_{\pi_{\rm DM}}+2n_{\eta}$, both equations combine giving $\dot{n}+3Hn=-\left(n_{\eta}n_{\pi_{\rm DM}}\langle\sigma_{\eta\pi}v\rangle-n_{\pi_{\rm DM}}^{2}\langle\sigma_{\pi\pi}v\rangle\right)$, where $\sigma_{\eta\pi}\equiv\sigma(\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM})$, and similarly for the inverse process. This can be simplified using detailed balance, $n^{\rm eq}_{\eta}\langle\sigma_{\eta\pi}v\rangle=n^{\rm eq}_{\pi_{\rm DM}}\langle\sigma_{\pi\pi}v\rangle$, as well as the fact that $\eta\leftrightarrow\pi_{\rm DM}\pi_{\rm DM}$ establish chemical equilibrium, $n_{\eta}/n_{\pi_{\rm DM}}^{2}=(n_{\eta}/n_{\pi_{\rm DM}}^{2})_{\rm eq}$, or in terms of chemical potentials, $\mu_{\eta}=2\mu_{\pi_{\rm DM}}$. Putting everything together, the yield for the combination $Y=Y_{\pi_{\rm DM}}+2Y_{\eta}$ satisfies
$\frac{dY}{dz}=-\langle\sigma_{\eta\pi}v\rangle\frac{sY_{\eta,\rm eq}}{zH}\left(\frac{Y_{\pi_{\rm DM}}^{3}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}}\right),$ (7)
where as usual $Y_{X}=n_{X}/s$, $z\equiv m_{\rm DM}/T$, $H$ is the Hubble rate and $s$ the entropy density of the Universe. The DM relic abundance is obtained upon numerical integration as $Y\simeq Y_{\pi_{\rm DM}}$. We show our results in the left panel of Fig 2 for BM1. These are independent of $v_{R}$ (if $v_{R}\lesssim 0.1$), but depend on $\delta$, or equivalently on $r_{ud}$ if the small dependence of the masses on $v_{R}$ and $\theta$ is neglected. For BM2, depending on $N_{f}$, due to the large number of degenerate pions, the relic density is reproduced in parameter regions in tension with the cluster bound discussed below. This suggests that the DM must lie in a small representation, which can be done breaking the mass degeneracy as in BM1. This is a generic feature of other color and flavor groups future .
Notice that the relic density is independent of $\theta$ as long as $\eta\leftrightarrow\pi_{\rm DM}\pi_{\rm DM}$ keep the two species in chemical equilibrium at least until the time at which $\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$ freezes out (see e.g. Frumkin:2021zng ). This implies a condition $\theta\geq\theta_{\rm min}(v_{R},m_{\rm DM},f_{\pi})$, estimated requiring that the thermally-average inverse decay rate overcomes the Hubble rate at freeze-out. We find $\theta_{\rm min}\sim 10^{-4}$ for both benchmarks. Notice also that the relic abundance is determined by $s$-wave 2-to-2 processes (see Table 1), in contrast to the SIMP scenario based on velocity suppressed 3-to-2 annihilations.
When $\theta<\theta_{\min}$ or $v_{R}\gg 1$, the number changing interactions in Eq. (4) lead to 3-to-2 annihilation competing with the WZW term Kamada:2017tsq . However, in our benchmark models, these reproduce the observed DM relic density only in parameter regions excluded by cluster observations, to be discussed below.
Astrophysical implications. The $\theta$ angle could also induce resonant scattering among DM particles in present-epoch halos, see Fig. 1. The corresponding Breit-Wigner formula for the non-relativistic cross section is Chu:2018fzy ; Chu:2019awd
$\!\!\!\!\!\!\!\!\sigma(v)$ $=$ $\sigma_{0}+\frac{128\pi\,}{m_{\rm DM}^{2}v_{R}^{2}}\frac{\Gamma^{2}}{m_{\rm DM}^{2}\left(v^{2}-v_{R}^{2}\right)^{2}+4\Gamma^{2}v^{2}/v_{R}^{2}}\,,$ (8)
which receives a resonant enhancement for DM relative velocities, $v$, close to the mass combination, $v_{R}$, given in Eq. (6). For reference, in SI units, $v\sim$ 50 km/s in small-scale galaxies while $v\sim$2000 km/s in clusters of galaxies. The 1-loop $\eta$ self-energy is included in $\sigma(v)$ as it impacts the scattering process Chu:2018fzy . The specific parameters for every benchmark are reported in Table 1. We note that for BM1 all DM today consists of $\pi^{0}$ due to the efficient conversion of $\pi^{\pm}$ into $\pi^{0}$ in the early universe Katz:2020ywn . Moreover, the production of $\pi^{\pm}$ from $\pi^{0}$ in astrophysical halos is kinematically forbidden as $\delta\gg v^{2}$ today (which holds unless $\delta\lesssim 10^{-5}$, or equivalently, $r_{ud}\sim 1$). For BM2, due to the $SU(n)$ symmetry, $\sigma_{0}$ coincides with that reported in Ref. Hochberg:2014kqa .
Clusters of galaxies constrain the self-scattering cross-section per unit mass below 0.5 cm2/g Harvey:2015hha ; Bondarenko:2017rfu ; Harvey:2018uwf ; Sagunski:2020spe ; DES:2023bzs . Regardless of $v_{R}$, the bound applies directly on $\sigma_{0}/m$, as depicted in Fig. 2 for each benchmark. We note that ordinary SIMPs obtaining their relic density from annihilations induced by the WZW term are in tension with this bound, as this happens in parameter regions in which $\sigma_{0}/m_{\rm DM}\gtrsim 1$ cm2/g (assuming chiral perturbation theory holds, i.e $m_{\rm DM}/f_{\pi}>4\pi/\sqrt{N_{c}}$). The fact that the cross section is independent of the velocity when $\theta=0$ aggravates this tension.
[…]
In this context, Ref. Chu:2018fzy pointed out that the velocity dependence encoded in Eq. (8) can successfully account for the required interactions in small-scale halos while evading the constraints arising from cluster observations. Following this study, we illustrate in Fig. 2 (right panel) the velocity-average cross section at different scales for three different values of $v_{R}$. The parameters are selected such that the mechanism detailed in the preceding section yields the observed DM density. Furthermore, they are also chosen so that $\sigma(v)/m_{\rm DM}\lesssim 100\,\text{cm}^{2}/\text{g}$ at $v\lesssim 100\,$km/s, due to non-observation of gravothermal collapse in dwarf-sized halos Tulin:2017ara ; Adhikari:2022sbh . This is rather conservative as the observational implications of gravothermal collapse for resonant SIDM are still under investigation Tran:2024vxy ; Kamada:2023dse (see also Refs. Yang:2022hkm ; Yang:2022zkd ; Outmezguine:2022bhq ). Interestingly, the behavior in Fig. 2 is favored by observations of ultra-faint dwarf galaxies, which suggest a sharp velocity dependence Hayashi:2020syu .
Although in this letter we remain agnostic about the origin of the required values of $v_{R}$ by the SIDM hypothesis –of order $10^{-4}$ in natural units– we highlight that the $\theta$ angle itself could account for their smallness. For instance, any mechanism enforcing $\mu=5m$ in BM2 leads to $v_{R}=0$ for $\theta=0$ and hence to $v_{R}\sim\theta$, as explained above. Furthermore, resonances with similar $v_{R}$ are found in nature, for example, in resonant collisions of $\alpha$-particles mediated by ${}^{8}\mathrm{Be}$, a very important process in stars, where $\alpha\alpha\rightarrow{}^{8}\mathrm{Be}$ followed by ${}^{8}\mathrm{Be}\,\alpha\rightarrow{}^{12}\mathrm{C}^{*}$ takes place, in analogy to the process setting the relic density above. For other examples in QCD-like models of DM, see Ref. Tsai:2020vpi .
Another signature of these scenarios is the emission of gravitational waves (GW) if the chiral PT is first order, see e.g. Refs. Witten:1984rs ; Schwaller:2015tja ; Helmboldt:2019pan ; Reichert:2021cvs . While in ordinary QCD the chiral PT is a smooth crossover Aoki:2006we ; Bhattacharya:2014ara , studies of dark QCD sectors with three or more light fermions, $m_{q}\ll\,\Lambda$, establish that the PT may be first order Schwaller:2015tja , with a critical temperature $T_{*}\sim f_{\pi}$, potentially allowing for GW emission. For the specific case of $\theta=0$, utilizing several effective models, Ref. Helmboldt:2019pan presented QCD-like benchmarks for critical temperatures of order $T_{*}~{}\sim 100$ MeV with a GW spectrum peaking at frequencies in the sub-mHz band (see also Ref. Reichert:2021cvs ). On the other hand, a non-vanishing $\theta$ may affect the PT: even with the matter content of SM QCD, $\theta\sim\pi$ could render the PT first order Bai:2023cqj , with potential consequences on the GW spectrum. This together with the fact that the dynamics of the PT and the GW spectrum depend on parameters and particle interactions not necessarily related with the pseudo-Goldstone bosons motivates a dedicated study for our benchmarks to draw conclusive statements. This is particularly important in light of the nHz signal from Pulsar Timing Arrays confirmed by the NANOGrav collaboration EPTA:2023fyk ; Reardon:2023gzh ; NANOGrav:2023gor ; Xu:2023wog , which could be explained with a PT with $T_{*}\sim 10-100$ MeV NANOGrav:2023hvm , in precise alignment with the viable values of $f_{\pi}$ in Fig. 2. See Ref. Han:2023olf for a hypothetical connection between the NANOGrav signal and SIDM with light mediators. The fact that our SIDM scenario does not require a light mediator could potentially alleviate the tension between this interpretation of the signal and cosmological observables requiring all unstable dark sectors particles to decay into the SM before $T\sim 1$ MeV Bringmann:2023opz .
Summary and outlook. We pointed out a production mechanism of DM in QCD-like theories where $\theta\neq 0$. This requires a resonance mediating $\pi_{\rm DM}\pi_{\rm DM}\to\eta$ followed by $\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$, with a mass such that $v_{R}\lesssim 0.1$. Several resonances of this type are known to exist in the SM, particularly in the QCD sector Tsai:2020vpi . Although it shares some similarities with theories where the relic density is obtained from annihilations induced by WZW term, here the viable parameter regions can be reconciled with cluster observations on DM self-scattering and are possibly far from the breakdown of chiral perturbation theory, see Fig. 2. This mechanism works even for small values of the topological angle, $\theta\gtrsim 10^{-4}$.
[…]
The $\theta$-dependent effects discussed here are generic features of QCD-like theories. For instance, in analogy with SM, Eq. (1) exhibits an anomalous axial $U(1)$ symmetry, whose pseudo-Goldstone boson, $\eta^{\prime}$, has $\theta$-dependent cubic vertices Witten:1980sp ; DiVecchia:1980yfw , potentially allowing for resonant scattering. Beyond $SU(N_{c})$ gauge groups, qualitatively similar effects arise for $SO(N_{c})$ or $Sp(N_{c})$. In particular, with $N_{c}\geq 4$, gauge confinement breaks flavor symmetry as $SU(N_{f})\to SO(N_{f})$ for $SO(N_{c})$ or $SU(N_{f})\to Sp(N_{f})$ for $Sp(N_{c})$. This leads to Lagrangian interactions analogous to Eq. (4) if $N_{f}\geq 3$ for $SO(N_{c})$ or $N_{f}\geq 6$ for $Sp(N_{c})$ (see the Appendix and Refs. Sannino:2016sfx ; LANDINIGIACOMO2022DMag ; Buttazzo:2019mvl ; Witten:1982fp ). Models of this kind will be presented in an upcoming publication future , together with an analysis of the Boltzmann equations.
Additionally, to be explored in future investigations are the CP violation effects induced by $\theta$-angle. Following Refs. Cline:2017qpe ; Carena:2018cjh it would be possible to leverage the novel source of CP violation in QCD-like sectors with a $\theta$-angle to generate a matter-antimatter asymmetry. This approach would offer a potential workaround to the tight constraints imposed by searches for electric dipole moments in scenarios with CP-violation in the visible sector.
Acknowledgments. We thank Xiaoyong Chu, Pilar Hernández, Manoj Kaplinghat, Hyungjin Kim and Tomer Volansky for useful discussions and Ayuki Kamada for comments on the manuscript. C.G.C. is supported by a Ramón y Cajal contract with Ref. RYC2020-029248-I, the Spanish National Grant PID2022-137268NA-C55 and Generalitat Valenciana through the grant CIPROM/22/69. G.L. is supported by the Generalitat Valenciana APOSTD/2023 Grant No. CIAPOS/2022/193. G.L. also acknowledges the hospitality of Universidade de Coimbra, Portugal. O.Z. has been partially supported by Sostenibilidad-UdeA, the UdeA/CODI Grants 2022-52380 and 2023-59130, and Ministerio de Ciencias Grant CD 82315 CT ICETEX 2021-1080. O.Z. also acknowledges the support of the Simons Foundation under Award number 1023171-RC and the hospitality of the International Institute of Physics at the Universidade Federal do Rio Grande do Norte in Natal, Brazil.
Appendix A Chiral Lagrangians in a $\theta$-vacuum
In the main text we illustrate the effects of a non-vanishing $\theta$ angle in a dark $SU(N_{c})$ QCD-like theory with $N_{f}$ flavors of light fermions in the fundamental representation and $N_{c}\geq 3$. Our results can be easily generalized to other choices of the gauge group, namely $SO(N_{c})$ or $Sp(N_{c})$. We first summarize the $SU(N_{c})$ case and then discuss the other possibilities.
Unitary groups. The Lagrangian describing $N_{f}$ massless Dirac fermions in the fundamental representation of a $SU(N_{c})$ gauge group with $N_{c}\geq 3$ enjoys a global flavor symmetry, $G_{F}=SU(N_{f})_{L}\times SU(N_{f})_{R}$, under which the fermion chiral components transform as $q_{L(R)}\to\exp{(i\alpha_{L(R)}^{a}\lambda^{a})}q_{L(R)}$, where $\lambda^{a}$ are the generators of $SU(N_{f})$ normalized as ${\rm Tr}(\lambda^{a}\lambda^{b})=2\delta^{ab}$. Confinement of gauge interactions induces a fermion condensate $\langle\bar{q}q\rangle\sim\Lambda^{3}$ which spontaneously breaks the flavor group $SU(N_{f})_{L}\otimes SU(N_{f})_{R}\to SU(N_{f})_{V}$. The unbroken vectorial subgroup is defined as the set of transformations for which $\alpha^{a}_{L}=\alpha^{a}_{R}$. This gives rise to $N_{f}^{2}-1$ Goldstone bosons, $\pi^{a}$, living in the coset space $(SU(N_{f})_{L}\otimes SU(N_{f})_{R})/SU(N_{f})_{V}\sim SU(N_{f})$. Their interactions are described by the effective Lagrangian $\mathcal{L}_{\rm eff}^{M=0}={f_{\pi}^{2}}{\rm Tr}[\partial_{\mu}U^{\dagger}\partial^{\mu}U]/4$, where $U=e^{i\Pi/f_{\pi}}$, with $\Pi=\pi^{a}\lambda^{a}$. A quark mass matrix $M$ explicitly breaks the flavor symmetry group $G_{F}$. However, if $m_{q}\ll\Lambda$, $G_{F}$ is still a good approximate symmetry, in which case, the small breaking due to the mass matrix is parametrized in the effective Lagrangian as
$\mathcal{L}_{\rm eff}=\frac{f_{\pi}^{2}}{4}{\rm Tr}[\partial_{\mu}U^{\dagger}\partial^{\mu}U]+\frac{f_{\pi}^{2}}{2}B_{0}{\rm Tr}[M_{\theta}U+U^{\dagger}M_{\theta}^{\dagger}]\,.$ (9)
Here, we include the effect of a non-vanishing $\theta$ term as $M_{\theta}=e^{i\theta Q/2}Me^{i\theta Q/2}$, with $Q=M^{-1}/{\rm Tr}M^{-1}$. The $\pi^{a}$ acquire a small mass $m_{\pi}\sim\sqrt{m_{q}\Lambda}\ll\Lambda$, becoming pseudo-Goldstone bosons.
The presence of $\theta\neq 0$ generates interactions involving an odd number of mesons. The leading terms are
$\mathcal{L}_{\theta}=\frac{\theta}{6f_{\pi}{\rm Tr}M^{-1}}B_{0}{\rm Tr}\Pi^{3}-\frac{\theta}{120f_{\pi}^{3}{\rm Tr}M^{-1}}B_{0}{\rm Tr}\Pi^{5},$ (10)
which gives rise to decays $\pi^{a}\to\pi^{b}\pi^{c}$, as well as 3-to-2 processes $\pi^{a}\pi^{b}\pi^{c}\to\pi^{d}\pi^{e}$. We introduce the $SU(N)$ structure constants: the totally anti-symmetric $f_{abc}=-i{\rm Tr}([\lambda_{a},\lambda_{b}]\lambda_{c})/4$ and the totally symmetric $d_{abc}={\rm Tr}(\{\lambda_{a},\lambda_{b}\}\lambda_{c})/4$. Using the $SU(N)$ identity ${\rm Tr}(\lambda_{a}\lambda_{b}\lambda_{c})=2(d_{abc}+if_{abc})$ and the symmetry properties of the structure constants we get
$\!\!\!\mathcal{L}_{\theta}=\frac{B_{0}\theta}{3f_{\pi}{\rm Tr}M^{-1}}\bigg{(}d_{abc}\pi_{a}\pi_{b}\pi_{c}-\frac{c_{abcde}}{10f_{\pi}^{2}}\pi_{a}\pi_{b}\pi_{c}\pi_{d}\pi_{e}\bigg{)},$ (11)
where $c_{abcde}=(\delta_{ab}d_{cde}+\delta_{cd}d_{abe})/N_{f}+\frac{1}{2}d_{abm}d_{cdn}d_{mne}$. These interactions are present whenever the $d_{abc}$ tensor is non-vanishing, which occurs if $N_{f}\geq 3$.
Finally, the Lagrangian must be augmented with the well-known Wess-Zumino-Witten (WZW) term, describing the 5-point interaction
${\cal L}_{\rm WZW}=-\frac{N_{c}}{240\pi^{2}f_{\pi}^{5}}\epsilon^{\mu\nu\rho\sigma}\operatorname{Tr}\left[\Pi\partial_{\mu}\Pi\partial_{\nu}\Pi\partial_{\rho}\Pi\partial_{\sigma}\Pi\right]\,.$ (12)
The WZW term induces an extra contribution to 3-to-2 processes $\pi^{a}\pi^{b}\pi^{c}\to\pi^{d}\pi^{e}$, which competes with that of $\theta$ and is non-vanishing if the coset space has a non-trivial fifth homotopy group, $\pi_{5}$. For unitary gauge groups, this occurs when $N_{f}\geq 3$, as in such a case $\pi_{5}(SU(N_{f}))=\mathbb{Z}$.
The case $N_{c}=2$ is special as the flavor group is $SU(2N_{f})$, whose corresponding fermion condensate breaks to $Sp(2N_{f})$, giving rise to $12N_{f}^{2}-N_{f}-$ 1 pseudo-Goldstone bosons. Similar to the case of $Sp(N_{c})$ discussed below, for $N_{f}\geq 3$ the $d_{abc}$ tensors do not vanish, while WZW term appears already for $N_{f}\geq 2$.
Orthogonal groups. The Lagrangian of $N_{f}$ massless Weyl fermions in the fundamental representation of $SO(N_{c})$ enjoys a $SU(N_{f})$ flavor symmetry. As usual, we consider $N_{c}\geq 4$ because $N_{c}=2$ and $N_{c}=3$ respectively correspond to $U(1)$ and $SU(2)$. The fermion condensate $\langle qq\rangle$ spontaneously breaks the flavor symmetry as $SU(N_{f})\to SO(N_{f})$, leading to $N_{f}(N_{f}+1)/2-1$ pseudo-Goldstone bosons. Their dynamics is described again by Eq. (9) with the $\lambda^{a}$ matrices in the $U$ field replaced with the elements of the coset $SU(N_{f})/SO(N_{f})$, which are the real generators of $SU(N_{f})$. Note that the real generators of $SU(N_{f})$ are symmetric while the imaginary ones are antisymmetric and correspond to those of $SO(N_{f})$. The odd interactions in Eq. (11) are present if the symmetric $d_{abc}$ tensor is non-vanishing, which occurs if $N_{f}\geq 3$. The WZW term is non-vanishing also for $N_{f}\geq 3$ as in that case $\pi_{5}(SU(N_{f})/SO(N_{f}))=\mathbb{Z}$ Hochberg:2014kqa ; Sannino:2016sfx .
Symplectic groups. The Lagrangian of $N_{f}$ massless Weyl fermions in the fundamental representation of a $Sp(N_{c})$ gauge group enjoys a $SU(N_{f})$ flavor symmetry, which is broken as $SU(N_{f})\to Sp(N_{f})$ by the fermion condensate. Notice that $N_{c}$ needs to be an even integer for symplectic groups, and without loss of generality, $N_{c}\geq 4$ as $Sp(2)=SU(2)$. Furthermore, $N_{f}$ must be even to avoid the global anomaly of the gauge group Witten:1982fp . The number of pseudo-Goldstone bosons is $N_{f}(N_{f}-1)/2-1$. The interactions among the mesons are described again by Eq. (9) with the meson field replaced by $U=e^{i\pi^{a}\lambda^{a}/f_{\pi}}\gamma_{N_{f}}$, where $\gamma_{N_{f}}\equiv\textbf{1}_{N_{f}/2}\otimes i\sigma_{2}$ is the antisymmetric invariant tensor of $Sp(N_{f})$. Here, the $\lambda^{a}$ matrices are the elements of the coset $SU(N_{f})/Sp(N_{f})$ which can be expressed as $\{\tilde{\lambda}_{R}^{a}\otimes\textbf{1}_{2},\tilde{\lambda}_{I}^{a}\otimes\sigma_{k}\}/\sqrt{2}$, where $\tilde{\lambda}_{R(I)}^{a}$ are the real symmetric (imaginary anti-symmetric) generators of $SU(N_{f}/2)$ and $\sigma_{k}$ are the Pauli matrices, see e.g. Refs. LANDINIGIACOMO2022DMag ; Buttazzo:2019mvl . The $d_{abc}$ tensor entering the odd interactions is non-vanishing for $N_{f}\geq 6$, see e.g. Ref. Kamada:2017tsq , while the WZW term is non-vanishing for $N_{f}\geq 4$ as $\pi_{5}(SU(N_{f})/Sp(N_{f}))=\mathbb{Z}$ in that case Hochberg:2014kqa .
Appendix B Explicit Lagrangian for each benchmark
Tadpoles. The interactions described in the chiral Lagrangian of Eq. (9) induce tadpoles, i.e. linear terms in the meson field $\Pi=\pi^{a}\lambda^{a}$. At leading order in $\theta$, these arise in the form ${\rm Tr}[\Pi\{M,Q\}]$. Choosing $Q=M^{-1}/{\rm Tr}M^{-1}$, which satisfies ${\rm Tr}Q=1$, and using ${\rm Tr}\Pi=0$, the linear terms vanish and Lagrangian takes the form of Eq. (3) of the main text up to ${\cal O}(\theta^{2})$ corrections. At higher orders in $\theta$, one must choose a more involved function $Q(\theta)$, which reduces to the expression above for $\theta\ll 1$.
Mass spectrum. The mass spectrum of the mesons is obtained expanding at the quadratic order in $\Pi$, which gives $\mathcal{L}_{\rm mass}=-B_{0}\pi^{a}\pi^{b}{\rm Tr}[M\{\lambda^{a},\lambda^{b}\}]/4+{\cal O}(\theta^{2})$. In the following we present the details for the two benchmark models discussed in the main text.
  • Benchmark BM1. We fix $N_{f}=3$ and we adopt the same notation of ordinary QCD. The quark mass matrix is $M={\rm diag}(m_{u},m_{d},m_{s})$ with $m_{u}\leq m_{d}\leq m_{s}$. The spectrum consists of 8 pseudo-Goldstone bosons: $\pi^{\pm}=(\pi_{1}\pm i\pi_{2})/\sqrt{2}$, $K^{\pm}=(\pi_{4}\pm i\pi_{5})/\sqrt{2}$, $K^{0}/\bar{K}^{0}=(\pi_{6}\pm i\pi_{7})/\sqrt{2}$ and Scherer:2002tk
    $\bigg{(}\begin{array}[]{c}\pi^{0}\\ \eta\end{array}\bigg{)}=\bigg{(}\begin{array}[]{cc}\cos{\theta_{\eta\pi}}&\sin{\theta_{\eta\pi}}\\ -\sin{\theta_{\eta\pi}}&\cos{\theta_{\eta\pi}}\end{array}\bigg{)}\bigg{(}\begin{array}[]{c}\pi_{3}\\ \pi_{8}\end{array}\bigg{)},\qquad\text{ with }\qquad\tan(2\theta_{\eta\pi})=\frac{\sqrt{3}(m_{u}-m_{d})}{(m_{u}+m_{d}-2m_{s})}.$ (13)
    The masses squared of the mesons are $m^{2}_{\pi^{\pm}}=B_{0}(m_{u}+m_{d})$, $m^{2}_{K^{\pm}}=B_{0}(m_{u}+m_{s})$, $m^{2}_{K,\bar{K}^{0}}=B_{0}(m_{d}+m_{s})$, while $m^{2}_{\pi^{0}}$ and $m_{\eta}^{2}$ are the eigenvalues of
    $\mathcal{M}^{2}_{\pi^{0},\eta}=\bigg{(}\begin{array}[]{cc}B_{0}(m_{u}+m_{d})&{B_{0}}(m_{u}-m_{d})/{\sqrt{3}}\\ {B_{0}}/(m_{u}-m_{d})/{\sqrt{3}}&{B_{0}}(m_{u}+m_{d}+4m_{s})/3\end{array}\bigg{)}.$ (14)
    The parameter space of interest in this work is that of resonant scattering, $m_{\eta}=(2+v_{R}^{2}/4)m_{\pi^{0}}$ with $v_{R}\lesssim 0.1$ constraining the full spectrum: up to a negligible dependence on $v_{R}$, the ratios among the meson masses are fixed in term of the parameter $r_{ud}=m_{u}/m_{d}$. In particular, this implies that $m_{\pi^{\pm}}=(1+\delta)m_{\pi^{0}}$, with $\delta$ varying from 0 to 0.075, as Fig. 3 shows.
  • Benchmark BM2. We fix $N_{f}=n+1$ with $n\geq 3$ and choose the quark mass matrix as $M={\rm diag}(m,..m,\mu)$ with $0<m<\mu$. There is a remnant $SU(n)$ symmetry under which the $N_{f}^{2}-1$ mesons organize as $n^{2}-1$ in the adjoint of $SU(n)$ –the pions $\pi$– as well as $2n$ in the (anti)fundamental representations –the kaons $K$–and one singlet corresponding to the $\eta$ resonance. Their masses squared are respectively $m^{2}_{\pi}=2B_{0}m$, $m^{2}_{K}=B_{0}(m+\mu)$ and $m^{2}_{\eta}=2B_{0}\left(m+n\mu\right)/(n+1)$. The expression $m_{\eta}=(2+v_{R}^{2}/4)m_{\pi}$ translates to $\mu=m[(n+1)(2+v_{R}^{2}/4)^{2}-1]/n$, which, for $v_{R}\lesssim 0.1$, implies $m_{K}^{2}=B_{0}m(5+3/n)+\mathcal{O}(v_{R}^{2})$. For $N_{f}=4$ this corresponds to $\mu\simeq 5m$, $m_{\eta}\simeq 2m_{\pi}$, $m_{K}\simeq\sqrt{3}m_{\pi}$.
Figure 3: Left: Splitting of the masses of the pions $\delta=m_{\pi^{\pm}}/m_{\pi^{0}}-1$ as a function of $r_{ud}$. Right: Masses of the 8 mesons normalized to $m_{\pi^{0}}$ as a function of $r_{ud}$. In both plots $v_{R}\lesssim 0.1$.
Cubic interactions induced by $\theta$. The interactions in Eq. (11) predict the decay of the $\eta$ meson. The relevant structure constants are $d_{888}=-d_{338}=-1/\sqrt{3}$ in the BM1 model, while $d_{\pi_{i}\pi_{i}\eta}=\sqrt{2/n(n+1)}$, with $i=1,\dots,n^{2}-1$, in the BM2 model. The corresponding cubic terms are
$\mathcal{L}^{(\rm BM1)}_{\eta\pi\pi}=\frac{B_{0}\theta}{\sqrt{3}f_{\pi}{\rm Tr}M^{-1}}\cos(3\theta_{\eta\pi})\eta\pi^{0}\pi^{0},\qquad\mathcal{L}^{(\rm BM2)}_{\eta\pi\pi}=\frac{B_{0}\theta}{\sqrt{n(n+1)/2}f_{\pi}{\rm Tr}M^{-1}}\,\eta\,\pi\cdot\pi.$ (15)
Notice that the expression for BM2 reduces to the one in BM1 if $n=2$ and for vanishing mixing, $\theta_{\eta\pi}=0$.
Appendix C Boltzmann equations and Dark Matter relic density
In this section we describe the computation of the DM relic density in more detail. We focus on the parameter space for which $v_{R}\lesssim 0.1$. In this regime, our results are independent on the precise value of $v_{R}$. All the stable mesons contribute to the DM relic density, although their contribution depends on their mass, as in usual scenarios of co-annihilating DM, where the contribution of the heavier particles is exponentially suppressed.
Derivation of the Boltzmann equations. In general, the spectrum can be grouped in the DM particle and the other stable mesons –collectively referred to as $\pi_{\rm DM}$– and the $\eta$ resonance, which decays via $\theta$-induced interactions. As usual, we introduce the parameter $z=m_{\rm DM}/T$ and the yield as $Y_{X}=n_{X}/s$, where $n_{X}$ is the number density and $s=2\pi^{2}g_{s}T^{3}/45$ is the total entropy density of the Universe. The Hubble rate in a radiation-dominated Universe is $H\simeq 1.66\sqrt{g_{*}}T^{2}/M_{\rm Pl}$ ( $g_{*}$ being the number of relativistic d.o.f.). All the stable mesons convert to DM by means of processes that maintain chemical equilibrium among the different species. This allows us to write a Boltzmann equation for the sum of the yields of DM and its co-annihilating partners, $Y_{\pi_{\rm DM}}$. Then, the evolution of $\pi_{\rm DM}$ and $\eta$ populations is described by the following coupled Boltzmann equations:
$\!\!\!\!\!\!\!\begin{cases}\begin{split}sHz\frac{dY_{\pi_{\rm DM}}}{dz}=&-\gamma_{3}({\pi_{\rm DM}\pi_{\rm DM}\pi_{\rm DM}}\to{\pi_{\rm DM}\pi_{\rm DM}})\left(\frac{Y_{\pi_{\rm DM}}^{3}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{3}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}\right)\\ &\!\!+2\gamma_{D}(\eta\to{\pi_{\rm DM}}{\pi_{\rm DM}})\left(\frac{Y_{\eta}}{Y_{\eta,{\rm eq}}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}\right)+\gamma_{2}(\eta\pi_{\rm DM}\to{\pi_{\rm DM}}{\pi_{\rm DM}})\left(\frac{Y_{\eta}}{Y_{\eta,{\rm eq}}}\frac{Y_{\pi_{\rm DM}}}{Y_{{\pi_{\rm DM}},{\rm eq}}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}\right)\end{split}\\ \begin{split}sHz\frac{dY_{\eta}}{dz}=&&-\gamma_{D}(\eta\to{\pi_{\rm DM}}{\pi_{\rm DM}})\left(\frac{Y_{\eta}}{Y_{\eta,{\rm eq}}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}\right)-\gamma_{2}(\eta{\pi_{\rm DM}}\to{\pi_{\rm DM}}{\pi_{\rm DM}})\left(\frac{Y_{\eta}}{Y_{\eta,{\rm eq}}}\frac{Y_{\pi_{\rm DM}}}{Y_{{\pi_{\rm DM}},{\rm eq}}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}\right),\\ \end{split}\end{cases}$ (16)
where $\gamma(I)$ is the equilibrium interaction rate density for the process $I$, obtained by adding the individual contributions of the co-annihilating partners. For $2\to 2$ scatterings, the latter is defined as
$\gamma(12\to 34)=S_{12}S_{34}\int d^{3}p_{1}d^{3}p_{2}f_{1}^{\rm eq}f_{2}^{\rm eq}\int d^{3}p_{3}d^{3}p_{4}(2\pi)^{4}\delta^{4}(p_{1}+p_{2}-p_{3}-p_{4})|\mathcal{A}_{12\to 34}|^{2},$ (17)
where $S_{12}(S_{34})=1\,(1/2)$ if the initial (final) particles are different (identical), $f^{\rm eq}_{i}$ is the equilibrium distribution for the particle $i$ and $\mathcal{A}$ is the amplitude of the process. In the non-relativistic limit, $m_{1,2}\gg T$, this reduces to $\gamma(12\to 34){\simeq}S_{12}n_{1,{\rm eq}}n_{2,{\rm eq}}\langle\sigma_{12\to 34}v\rangle,$ in terms of the thermal averaged cross section $\langle\sigma_{12\to 34}v\rangle$ and the equilibrium number densities. Similar expressions hold for $n\to 2$ processes with $n\geq 1$. Detailed balance implies $\gamma_{2}(\pi_{\rm DM}\pi_{\rm DM}\to\eta\pi_{\rm DM})=\gamma_{2}(\eta{\pi_{\rm DM}}\to{\pi_{\rm DM}}{\pi_{\rm DM}})$ and similarly for the other processes. Additional number-changing interactions, such as 4-to-2 annihilations are assumed negligible. Note that might not be the case for keV DM Bernal:2015xba .
Furthermore, for $v_{R}$ sufficiently small, 3-to-2 annihilations are dominated by the on-shell exchange of $\eta$ (see Fig. 1 of the main text). To avoid double-counting, from the total 3-to-2 rate, one must subtract the resonant piece, associated with $\pi_{\rm DM}\pi_{\rm DM}\to\eta$ followed by $\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$, see e.g. Ref. Kolb:1979qa . Here, $\gamma_{3}(\pi_{\rm DM}\pi_{\rm DM}\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM})$ only refers to the non-resonant piece after subtraction, which is small. In the same region of the parameter space, the inverse decay processes $\pi_{\rm DM}\pi_{\rm DM}\to\eta$ are strongly enhanced. As a result decays $\eta\to\pi_{\rm DM}\pi_{\rm DM}$ and inverse decay keep the $\pi_{\rm DM}$ fields and $\eta$ in chemical equilibrium, even after all other DM number-changing interactions have frozen-out. In terms of chemical potentials, $\mu_{\eta}=2\mu_{\pi_{\rm DM}}$ which leads to
$\frac{Y_{\eta}}{Y_{\eta,{\rm eq}}}=\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}.$ (18)
This only works as long as the decays and inverse decays $\eta\leftrightarrow\pi_{\rm DM}\pi_{\rm DM}$ are both active, which places a lower bound $\theta\gtrsim\theta_{\rm min}$, as discussed in the main text and detailed further below. We underline that the presence of the resonance is crucial, as for $m_{\eta}\gg m_{\rm DM}$, the inverse decay rate gets suppressed by a factor $e^{-m_{\eta}/m_{\rm DM}}$, freezing out much before than the other DM number-changing processes, invalidating Eq. (18). All in all, the Boltzmann Eqs. (16), under the condition Eq. (18) simplify to
$\begin{cases}\begin{split}sHz\frac{dY_{\pi_{\rm DM}}}{dz}=&-\bigg{(}\gamma_{2}(\eta{\pi_{\rm DM}}\to{\pi_{\rm DM}}{\pi_{\rm DM}})+\gamma_{3}({\pi_{\rm DM}\pi_{\rm DM}\pi_{\rm DM}}\to{\pi_{\rm DM}\pi_{\rm DM}})\bigg{)}\left(\frac{Y_{\pi_{\rm DM}}^{3}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{3}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}\right)\end{split}\\ \begin{split}sHz\frac{dY_{\eta}}{dz}=&-\gamma_{2}(\eta{\pi_{\rm DM}}\to{\pi_{\rm DM}}{\pi_{\rm DM}})\left(\frac{Y_{\pi_{\rm DM}}^{3}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}\right)\\ \end{split}\end{cases}.$ (19)
As expected, the processes $\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$ change the number of DM particles as the 3-to-2 reactions do. We can then write a single Boltzmann equation for the combination $Y_{\pi_{\rm DM}}+2Y_{\eta}$,
$\begin{split}sHz\frac{d(Y_{\pi_{\rm DM}}+2Y_{\eta})}{dz}\simeq&-\gamma_{2}(\eta{\pi_{\rm DM}}\to{\pi_{\rm DM}}{\pi_{\rm DM}})\left(\frac{Y_{\pi_{\rm DM}}^{3}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{3}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}\right),\end{split}$ (20)
where we neglect the non-resonant piece, $\gamma_{3}(\pi_{\rm DM}\pi_{\rm DM}\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM})$. Furthermore, as $Y_{\eta}\ll Y_{{\pi_{\rm DM}}}$, we can approximate $Y_{\pi_{\rm DM}}+2Y_{\eta}\simeq Y_{\pi_{\rm DM}}$. In the non-relativistic limit, $m_{\rm DM}\gg T$, we can write $\gamma_{2}\simeq s^{2}Y_{{\pi_{\rm DM}},{\rm eq}}Y_{\eta,{\rm eq}}\langle\sigma_{\eta\pi}v\rangle$, with $\sigma_{\eta\pi}\equiv\sigma(\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM})$, so that
$\frac{dY_{\pi_{\rm DM}}}{dz}\simeq-\langle\sigma_{\eta\pi}v\rangle\frac{sY_{\eta,\rm eq}}{zH}\left(\frac{Y_{\pi_{\rm DM}}^{3}}{Y_{{\pi_{\rm DM}},{\rm eq}}^{2}}-\frac{Y_{\pi_{\rm DM}}^{2}}{Y_{{\pi_{\rm DM}},{\rm eq}}}\right).$ (21)
The DM relic density is obtained upon numerical integration of the previous equation. As long as the condition $\theta\gtrsim\theta_{\rm min}$ is full-filled, the DM relic density is independent of the precise value of $\theta$. Our results are well approximated by evaluating the DM relic density as $Y_{\pi_{\rm DM},\rm eq}(z_{\rm fo})$ where $z_{\rm fo}$ is the time at which the scatterings $\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$ freeze out, estimated as $H(z_{\rm fo})\approx n_{\eta,{\rm eq}}(z_{\rm fo})\langle\sigma_{\eta\pi}v\rangle$. This gives (assuming constant $g_{s}\simeq g_{*}$)
$z_{\rm fo}\simeq\frac{m_{\rm DM}}{m_{\eta}}\log\left[\frac{\sqrt{z_{\rm fo}}}{1.66\sqrt{g_{*}}}\left(\frac{m_{\eta}}{2\pi m_{\rm DM}}\right)^{3/2}m_{\rm DM}M_{\rm Pl}\langle\sigma_{\eta\pi}v\rangle\right],$ (22)
which can be solved iteratively. We find $z_{\rm fo}\sim 15-25$.
Effective cross sections. The annihilation cross section $\langle\sigma_{\eta\pi}\rangle$ depends on the specific benchmark model.
  • Benchmark BM2. The total DM relic density is the sum of the contributions of pions $\pi$ and kaons $K$, which are maintained in chemical equilibrium through 2-to-2 scatterings, which freeze out after $z_{\rm fo}$. Nonetheless, kaons are heavier and their abundance is exponentially suppressed with respect to the one of the pions as $Y_{K}/Y_{\pi}\sim\exp{[-(m_{K}-m_{\rm DM})z/m_{\rm DM}]}$. In the region of the parameter space that we are considering here, $m_{K}\simeq\sqrt{3}m_{\pi}$ and the exponential suppression is strong (roughly a factor $\lesssim 10^{-3}$ for $z>10$). Kaons can thus be safely neglected and $Y_{\pi_{\rm DM}}\simeq Y_{\pi}$. The only cross section relevant to determine the DM relic density is $\sigma(\eta\pi\to\pi\pi)\equiv\sigma_{\eta\pi}$. As a result, $\langle\sigma_{\eta\pi}v\rangle=\sqrt{5}m_{\rm DM}^{2}(n^{2}-4)/(192\pi f_{\pi}^{4}n^{2}(n+1))$.
  • Benchmark BM1. Here the spectrum is more involved: four kaons and $\pi^{\pm}$, which are stable, in addition to the lightest $\pi^{0}$ and the $\eta$ resonance. All the stable mesons are maintained in chemical equilibrium by conversion processes such as $\pi^{+}\pi^{-}\leftrightarrow\pi^{0}\pi^{0}$ (and similar for the other species), which freeze-out after $z_{\rm fo}$. While the contribution of kaons is irrelevant in analogy to the BM2 model, $\pi^{\pm}$ cannot be neglected, as $\delta\equiv m_{\pi^{\pm}}/m_{\pi^{0}}-1\lesssim 0.075$. In fact, at the freeze-out of $\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$, $\pi^{\pm}$ are almost as abundant as $\pi^{0}$. Therefore, $Y_{\pi_{\rm DM}}\simeq Y_{\pi^{0}}+Y_{\pi_{1}}+Y_{\pi_{2}}=Y_{\pi^{0}}+2Y_{\pi^{+}}$. We must take into account the $\eta\pi\to\pi\pi$ processes which involve all the 3 pions. At the leading order in the $\delta$ parameter these are given by
    $\langle\sigma(\eta\pi^{0}\to\pi_{1,2}\pi_{1,2})v\rangle=\frac{529\sqrt{5}\,m_{\pi^{0}}^{2}}{5184\pi f_{\pi}^{4}}\delta,$ $\langle\sigma(\eta\pi_{1,2}\to\pi^{0}\pi_{1,2})v\rangle=\frac{49\sqrt{5}\,\,m_{\pi^{0}}^{2}}{2592\pi f_{\pi}^{4}}\delta,\quad$ $\langle\sigma(\eta\pi^{0}\to\pi^{0}\pi^{0})v\rangle=\frac{\sqrt{5}\,m_{\pi^{0}}^{2}}{64\pi f_{\pi}^{4}}\delta,$ (23)
    which, after accounting for combinatoric factors, are summed as
    $\langle\sigma_{\eta\pi}v\rangle\simeq\frac{1}{3}\bigg{(}\langle\sigma(\eta\pi^{0}\to\pi^{0}\pi^{0})v\rangle+2\langle\sigma(\eta\pi^{0}\to\pi_{1}\pi_{1})v\rangle+2\langle\sigma(\eta\pi_{1}\to\pi^{0}\pi_{1})v\rangle\bigg{)}.$ (24)
    In the non-relativistic limit, this enters $\gamma(\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM})=n_{\eta,\rm eq}n_{\pi_{\rm DM},\rm eq}\langle\sigma_{\eta\pi}v\rangle$, to be replaced in Eq. (21).
    Notice that today DM is made only by $\pi^{0}$. The conversion processes $\pi^{+}\pi^{-}\to\pi^{0}\pi^{0}$ can efficiently deplete the $\pi^{\pm}$ population as long as the mass splitting is $\delta\gtrsim 10^{-5}$ Katz:2020ywn , which is realized in our setup as long as $r_{ud}\lesssim 0.95$.
Determination of $\theta_{\rm min}$
Figure 4: Left: Contours of $\theta_{\rm min}$ as a function of $m_{\rm DM}$ and $f_{\pi}$ in the BM1 model. Right: Thermal averaged interaction rates for the different processes compared to the Hubble rate, for values of the parameters that reproduce the DM relic abundance with $\theta>\theta_{\rm min}$. Notice that (inverse) decays keep $\eta$ and $\pi^{0}$ in chemical equilibrium even for $z>z_{\rm fo}$. Analogously $\pi^{+}\pi^{-}\leftrightarrow\pi^{0}\pi^{0}$ conversions keep the three pions in chemical equilibrium.
Eq. (18) works as long as $\pi_{\rm DM}$ and $\eta$ are kept in chemical equilibrium after the freeze-out of the reactions $\eta\pi_{\rm DM}\to\pi_{\rm DM}\pi_{\rm DM}$, occurring at $z_{\rm fo}$. This requires that $\theta$ is larger than a minimal value, which can be estimated as follows. We consider the thermal averaged decay and inverse decay rates ¡
$\langle\Gamma(\eta\to\pi_{\rm DM}\pi_{\rm DM})\rangle=\frac{K_{1}(m_{\eta}/T)}{K_{2}(m_{\eta}/T)}\Gamma(\eta\to\pi_{\rm DM}\pi_{\rm DM}),\qquad\langle\Gamma(\pi_{\rm DM}\pi_{\rm DM}\to\eta)\rangle=\frac{n_{\eta,{\rm eq}}}{n_{{\pi_{\rm DM}},{\rm eq}}}\langle\Gamma(\eta\to\pi_{\rm DM}\pi_{\rm DM})\rangle,$ (25)
where $K_{1,2}$ are the modified Bessel functions. These processes are active as long as
$\langle\Gamma(\eta\to\pi_{\rm DM}\pi_{\rm DM})\rangle>H\qquad\text{and}\qquad\langle\Gamma(\pi_{\rm DM}\pi_{\rm DM}\to\eta)\rangle>H.$ (26)
Since inverse decay rate is exponentially suppressed after $z\gtrsim 1$, to maintain the chemical equilibrium among $\eta$ and $\pi_{\rm DM}$ at $z\geq z_{\rm fo}$, it is necessary that
$\langle\Gamma(\pi_{\rm DM}\pi_{\rm DM}\to\eta)\rangle\geq H\qquad\text{at}\qquad z=z_{\rm fo}.$ (27)
This determines the minimal value $\theta_{\rm min}$ as a function of $(v_{R},m_{\rm DM},f_{\pi})$. We show some reference values in Fig. 4 (left) for the BM1 model. Similar results hold the BM2 model. Notice that the dependence on $v_{R}$ is not trivial: in the relevant region of the parameter space for this work, $v_{R}\lesssim 0.1$, we have $\Gamma(\pi_{\rm DM}\pi_{\rm DM}\to\eta)\propto v_{R}$. Hence, as $v_{R}$ gets smaller, larger values of $\theta_{\rm min}$ are required to compensate the suppression. In Fig. 4 (right) we show the interaction rates for the different relevant processes compared with the Hubble rate for different values of the parameters in the BM1 model. Clearly, the DM production mechanism studied here works even for small values of the $\theta$ parameter, in the range of $10^{-4}$.